Grade 11General Physics

A ball of mass 100 g is projected upwards from the ground with a velocity of 49m/s . At the same time another ball is dropped from a height of 98 m to fall freely along the same path as ball. After some time they collide and stick together and fall together. Find the time of flight

Shashank Raut
10 Years agoGrade 11

1 Answer

Riddhish Bhalodia
10 Years ago
FIrst task is we find at waht time t and place x from the bottom they collide
for ball 1, g = 9.8
x = 49t - 4.9t^2
and for ball 2
98 - x = 4.9t^2
solving the two equations we get
t = 2seconds
x = 78.4m

Now as they stick it’s an inelastic collision which can be solved by momentum conservation but we need their velocities first
ball 1 : v1 = 49 – 9.8*2 = 29.4
ball 2 v2 = -19.6

by momentum conservation
mv1 + mv2 = 2mv
v = (v1+v2)/2
v = 4.9m/s (upwards direction)

Now time to reach ground from this point onwards is
78.4 = -4.9t + 4.9t^2
t^2 - t -16 =0
solving we get t = 4.83seconds
so total time of flight is
2 + 4.83 = 6.83 seconds